OpenStudy (anonymous):

help!! A circle has a diameter of 10 yards. What is its area? Use 3.14 for pi A. 31.4 Sq yd B. 39.25 Sq Yd C. 62.8 Sq Yd D. 78.5 Sq Yd

OpenStudy (calculusxy):

Remember this formula to figuring out the area of a circle. r^2(pi)=radius squared times pi. Since you only have the diameter, you know that the radius is half of the diameter. So half of 10yards is 5yards. The radius is now 5yards. When you square 5, you get 25 because 5*5=25. Now you multiply 25*3.14=78.50 The area of the circle is 78.50

OpenStudy (calculusxy):

So the answer is D.

OpenStudy (calculusxy):

Do you get what I mean?

OpenStudy (anonymous):

Yes thank you ^^

OpenStudy (calculusxy):

No problem!